The physician orders dopamine IV drip at 4 mcg/kg/min for low blood pressure. The patient weighs 180lb. The medication is supplied in 300 mg/ 500 mL of 0.45% NS. Calculate the flow rate in milliliters per hour. ______________

(please show work and where all numbers come from)
Answer #1

1 lb = 2.2 kg 180*2.2 = 396 kgs

Physician orders 4mcg/kg so 396*4 = 1584 mcg

1000mcg = 1 mg so 1584 mcg = 1.584 mg so around 1.6 mg. We have to infuse 1.6mg per minute .

300mg = 500 ml., 300%500 = 0.6..so 1ml= 0.6mg

0.6 * 2.7 = 1.62 mg , so we have to infuse 2.7 ml per minute to infuse 1.6mg of dopamine.per minute.

2.7 * 60 = 162, so we have to set 162ml over 1 hour